OpenStudy (anonymous):

Let f(x) = 2x + 2. Solve f-1(x) when x = 4. 1 3 4 10

OpenStudy (anonymous):

is it c or a

OpenStudy (happyfeet8040):

I'd say A, but I'm not completely sure..:/

OpenStudy (anonymous):

f(x) = 2x + 2. Solve f-1(x) when x = 4 = 1, A

OpenStudy (anonymous):

@e.mccormick

OpenStudy (happyfeet8040):

So it's A?

OpenStudy (anonymous):

i think a

OpenStudy (e.mccormick):

First, did you find the inverse, and if so what was it?

OpenStudy (anonymous):

i got a

OpenStudy (anonymous):

Let f(x) = x2 - 81. Find f-1(x). �square root of the quantity x plus 81 9square root of x 1 over quantity x squared minus 81 x squared over 81i think its d

OpenStudy (anonymous):

@LunyMoony

OpenStudy (e.mccormick):

Is d the over one?

OpenStudy (lunymoony):

B

OpenStudy (anonymous):

for which one

OpenStudy (e.mccormick):

And how are you getting that Luny?

OpenStudy (lunymoony):

wait, what you mean by this"f-1(x)" ?

OpenStudy (lunymoony):

nevermind I know, sec

OpenStudy (e.mccormick):

\(f(x) = x^2 - 81\). Find \(f^{-1}(x)\) means: \(x=y^2-81\) solve for y.

OpenStudy (anonymous):

my brain is about to exploed

OpenStudy (lunymoony):

A, I believe

OpenStudy (anonymous):

for which question

OpenStudy (lunymoony):

oh, sry was looking at the wrong one, sec

OpenStudy (lunymoony):

I think its D to, your last question

OpenStudy (anonymous):

Let f(x) = twenty-four over the quantity of 3 x minus two. Find f(-2). -6 -3 3 6 is it d

OpenStudy (anonymous):

@LunyMoony

OpenStudy (anonymous):

Let f(x) = x2 - 81. Find f-1(x). �square root of the quantity x plus 81 9square root of x 1 over quantity x squared minus 81 x squared over 81

OpenStudy (anonymous):

is it d

OpenStudy (lunymoony):

@Popraci

OpenStudy (anonymous):

f(x) = 24/(3x-2) f(x) = 24/(3*(-2) -2) f(-2) = -3

OpenStudy (anonymous):

luny is that ur gf

OpenStudy (anonymous):

@LunyMoony is @Popraci ur gf

OpenStudy (lunymoony):

no, hes a guy and he is a teacher's assiestent

OpenStudy (anonymous):

Which function below has the largest slope? f(x) x f(x) -3 -10 -1 -7 1 -4 ������g(x) g(x) equals one-half x plus 3 h(x) Lance gains 1 level each time he plays his favorite video game. He started at level 10. f(x) g(x) h(x) The functions all have the same slope

OpenStudy (lunymoony):

is there a question ?